# Dynamic Programming: From Local Optimality to Global Optimality

### v2

@@ 5.2. Reducible optimal savings MDP
Now consider a modified version of the optimal saving example where both returns and labor income are bounded:
@@
with $10<\underline{\eta}<\overline{\eta}<$ 1 and $0\leqslant\underline{y}<\overline{y}<\infty$. For $w\in\mathbbm{R}_{+}$ and a Borel set $B\subseteq\mathbbm{R}_{+}$, the stochastic kernel $P_{\sigma}$ is:
@@
where $\psi$ and $\varphi$ have support on $[\underline{\eta},\overline{\eta}]$ and $[\underline{y},\overline{y}]$ respectively. In this case, we let $\psi$ and $\varphi$ be uniform distributions.
Figure 3. The law of motion of wealth given consumption $c=0$ with $\overline{\eta}=0.8$ and $\overline{y}=8$.
The next proposition shows that this optimal saving MDP is reducible.
@@
For any feasible policy $\sigma\in\Sigma$, $P_{\sigma}$ is reducible.
@@
Fix initial wealth $w_{0}\in\mathbbm{R}_{+}$. For any $t$-step transition, let $\alpha_{t}=w_{t}-\sigma(w_{t})$ be the savings at step $t$. Then for any $\sigma\in\Sigma$
@@
$w_{n}<\overline{\eta}w_{0}+\overline{y}\frac{1}{1-\overline{\eta}}<M\quad\forall n\in\mathbbm{N}$
Let $B=(M,\infty)$. Then $\pi(B)>0$, and $P^{n}_{\sigma}(w_{0},B)=0$ for all $n\in\mathbbm{N}$ as $w_{n}$ is bounded above by $M$ for all $n\in\mathbbm{N}$. ∎
The failure of irreducibility can be understood through the dynamics illustrated in Figure 3. When starting from any initial wealth $w\in\mathsf{W}$, the law of motion restricts the wealth process to converge towards a steady state. This convergence prevents the process from exploring the entire state space. The figure specifically demonstrates the limited range of wealth levels that can be reached from a given initial state when following the policy $\sigma(x)=0$.
Figure 4. $\hat{v}_{\sigma}$ and $\hat{\sigma}$ with $w=1$ and $w=50$ against the OPI solutions with $[\underline{\eta},\overline{\eta}]=[0.5,0.8]$ and $[\underline{y},\overline{y}]=[1,8]$.
(a) $w=1$
(b) $w=50$
@@
Figure 5. Expected returns (13) and value loss (14) over training episodes for the reducible optimal saving model.
@@
For reducible MDPs, Algorithm 1 no longer guarantees convergence to the optimal policy when optimized at a single initial state $w\in\mathsf{W}$. This limitation is clearly demonstrated in Figure 4, which shows significant discrepancies between the algorithm’s output and the OPI solution. Figure 5 confirms this limitation, showing that extended training episodes fail to improve the policy’s performance.
@@
Figure 4 also provides empirical support for Theorem 4.1. The policy network achieves near-optimal performance at lower wealth levels because these levels are reachable from the initial state under the learned policy $\hat{\sigma}$. However, the performance deteriorates for wealth levels that are not reachable from the initial state given the policy $\hat{\sigma}$. We note the reachable range of wealth level in the graph is smaller than in Figure 3 because $\hat{\sigma}(x)>0$ for all $x\in\mathsf{W}$.
@@ 6. Extensions and Future Work
Using MDP optimality results from Bäuerle and Rieder (2011) or Bertsekas (2022), it should be possible to extend our results to the case of unbounded rewards by replacing the ordinary supremum norm on $b\mathsf{X}$ with a weighted supremum norm. Also, while our results have focused on standard MDPs with constant discount factors, one useful variation of this model is MDPs with state-dependent discount factors, so that $\beta$ becomes a map from $\mathsf{X}$ to $\mathbbm{R}_{+}$ (see, e.g., Stachurski and Zhang (2021)). We conjecture that similar results will be available under suitable stability and irreducibility assumptions. The ideas are left for future research.
It seems likely that results similar to Theorem 3.3 will be valid for some continuous time MDPs, as well as at least some of the nonstandard discrete time dynamic programs discussed in Bertsekas (2022) and Sargent and Stachurski (2025). These topics are also left for future work.

### v3

@@ 5.2. A Reducible MDP
Now consider a modified version of the optimal saving MDP where both returns and labor income are bounded:
@@
with $10<\underline{\eta}<\overline{\eta}<$ 1 and $0\leqslant\underline{y}<\overline{y}<\infty$. For $w\in\mathbbm{R}_{+}$ and a Borel set $B\subseteq\mathbbm{R}_{+}$, the stochastic kernel $P_{\sigma}$ is
@@
where $\psi$ and $\varphi$ have support on $[\underline{\eta},\overline{\eta}]$ and $[\underline{y},\overline{y}]$ respectively. In this case, we let $\psi$ and $\varphi$ be uniform distributions. For the rest of this section, we always set $[\underline{\eta},\overline{\eta}]=[0.5,0.8]$ and $[\underline{y},\overline{y}]=[1,8]$. At the same time, we deliberately continue to take $\mathbbm{R}_{+}$ as the state space. This leads to a failure of irreducibility, as clarified in the next proposition.
@@
For any $\sigma\in\Sigma$, the transition kernel $P_{\sigma}$ is neither weakly irreducible nor open set irreducible.
@@
Fix $\sigma\in\Sigma$. In view of Lemma 4.3, we only need to show that $P_{\sigma}$ is not weakly irreducible. To this end, let $\pi$ be a distribution on $\mathsf{W}$ that assigns positive probability to open sets. Fix initial wealth $w_{0}\in\mathbbm{R}_{+}$. For any $t$-step transition, let $\alpha_{t}\coloneq w_{t}-\sigma(w_{t})$ be savings at step $t$. Then
@@
$w_{n}<\overline{\eta}w_{0}+\overline{y}\frac{1}{1-\overline{\eta}}<M\quad\forall n\in\mathbbm{N}.$
Let $B=(M,\infty)$. Then $\pi(B)>0$, and $P^{n}_{\sigma}(w_{0},B)=0$ for all $n\in\mathbbm{N}$. This shows that $P_{\sigma}$ is not weakly irreducible. ∎
Figure 3. The upper bound law of motion for wealth
The proof of Proposition 5.2 is partly illustrated in Figure 3. The figure shows the 45 degree line and an upper bound law of motion for wealth, obtained by setting consumption to zero and both shocks to their upper bound. States above the steady state are not reachable from states below the steady state.
Figure 4. $\hat{v}_{\sigma}$ and $\hat{\sigma}$ with $\bar{w}=1$ vs OPI solutions
@@
Figure 5. $\hat{v}_{\hat{\sigma}}(\bar{w})$ over training episode for the reducible optimal saving model at $\bar{w}=1$.
@@
For reducible MDPs, Algorithm 1 no longer guarantees convergence to the optimal policy when optimized at a single initial state $\bar{w}\in\mathsf{W}$. This limitation is clearly demonstrated in Figure 5, which shows significant discrepancies between the algorithm’s output and the OPI solution. Figure 5 confirms this limitation, showing that extended training episodes fail to improve the policy’s performance. [^3] The training episode in Figure 5 extends longer than that in Figure 2 because we implement early stopping with a patience parameter of 150 episodes, meaning training terminates after 150 consecutive episodes without improvement in the loss.
@@
Figure 5 provides empirical support for Lemma 4.1. The policy network achieves near-optimal performance at lower wealth levels because these levels are reachable from the initial state under the learned policy $\hat{\sigma}$. However, the performance deteriorates for wealth levels that are not reachable from the initial state given the policy $\hat{\sigma}$. (Reachable range of wealth level in the graph is smaller than in Figure 3 because $\hat{\sigma}(w)>0$ for all $w\in\mathsf{W}$.) We also include the result for $\bar{w}=50$ in Appendix C to show that the same phenomenon related to the reachable subset occurs for larger initial wealth levels.
These results confirm that, when irreducibility fails, the policy $\hat{\sigma}$ computed to maximize local optimality is not guaranteed to be optimal, and the performance of the policy is sensitive to the initial state and the reachable subset of the state space starting from the initial state.
5.3. Firm Entry
The next example provides an extension showing how the results in the paper can be used outside the MDP setting. This is significant because many recent examples of dynamic programs are not MDPs (see, e.g., Sargent and Stachurski (2025a)). The example concerns an optimal stopping problem with state-dependent discounting. We prove that, under the assumptions stated below, local optimality implies global optimality.
Specifically, we consider a stopping problem for a firm choosing if and when to enter a market. Stopping corresponds to entering the market and continuing corresponds to waiting until the next period and then deciding again. If the firm enters at time $t$ it receives one-off profit $\pi(X_{t})$, where $(X_{t})_{t\geqslant 0}$ is a Markov sequence on possibly unbounded interval $\mathsf{X}\subset\mathbbm{R}$ with transition kernel $Q$. If the firm continues then it pays a fixed cost $c$. We admit the possibility that discounting is not constant. (This allows for the fact that firms evaluate future profit opportunities based on cost-of-capital, which fluctuates over time.) Thus, the Bellman equation for the firm is
$v(x)=\max\left\{\pi(x),\;-c+\beta(x)\int v(x^{\prime})Q(x,\mathop{}\!\mathrm{d}x^{\prime})\right\},$ (15)
where $\beta(x)\in(0,\infty)$ is the discount factor associated with cost-of-capital in state $x$. (Discount factor realizations above one are admitted in order to accommodate occasionally negative interest rates.)
Let $ib\mathsf{X}$ be the increasing functions in $b\mathsf{X}$ and let $icb\mathsf{X}$ be the continuous functions in $ib\mathsf{X}$. We assume that the discount operator $K$ defined by
$(Kf)(x)=\beta(x)\int f(x^{\prime})Q(x,\mathop{}\!\mathrm{d}x^{\prime})\qquad(x\in\mathsf{X},\;f\in b\mathsf{X})$
preserves monotonicity and continuity, in the sense that $K$ is invariant on both $ib\mathsf{X}$ and $icb\mathsf{X}$, and that $r(K)<1$, where $r(K)$ is the spectral radius of $K$. For simplicity, we also assume that $Q$ always has full support on $\mathbbm{R}$, so that $Q(x,B)>0$ for any Borel set $B\subset\mathbbm{R}$ with positive Lebesgue measure. (For example, $(X_{t})$ might be driven by an AR(1) sequence with normal shocks.) Finally, we assume that the profit function $\pi$ is bounded, increasing, and continuous.
A policy is a map $\sigma$ from $\mathsf{X}$ to $\{0,1\}$, with $\sigma(x)=1$ indicating the decision to stop. The lifetime value $v_{\sigma}$ of policy $\sigma$ is the unique fixed point of the operator $T_{\sigma}$ on $b\mathsf{X}$ given by
$(T_{\sigma}\,v)(x)=\sigma(x)\pi(x)+(1-\sigma(x))\left[-c+\beta(x)\int v(x^{\prime})Q(x,\mathop{}\!\mathrm{d}x^{\prime})\right].$ (16)
We let $\Sigma$ be the set of all policies $\sigma$ such that $T_{\sigma}$ is invariant on $ib\mathsf{X}$.
As for the MDP case, the value function is denoted $v^{*}$ and defined at each $x\in\mathsf{X}$ by $v^{*}(x)\coloneq\sup_{\sigma\in\Sigma}v_{\sigma}(x)$. A policy $\sigma$ is called optimal if $v_{\sigma}(x)=v^{*}(x)$ for all $x\in\mathsf{X}$. In the proofs below we will make use of the Bellman operator for this problem, which takes the form
$(Tv)(x)=\max\left\{\pi(x),\;-c+\beta(x)\int v(x^{\prime})Q(x,\mathop{}\!\mathrm{d}x^{\prime})\right\}\qquad(x\in\mathsf{X},\;v\in b\mathsf{X}).$
The model has standard optimality properties, which we detail in the theorem below.
Theorem 5.4.
Under the stated assumptions, the following results hold:
  1. (a)
    The value function is the unique solution to the Bellman equation in $b\mathsf{X}$,
  2. (b)
    the value function is increasing and continuous, and
  3. (c)
    at least one optimal policy exists.
We will prove in addition the following result, which holds for any $\sigma\in\Sigma$:
Theorem 5.5.
If there exists an $x\in\mathsf{X}$ such that $v_{\sigma}(x)=v^{*}(x)$, then $\sigma$ is optimal.
We begin the proof of Theorem 5.4 and Theorem 5.5 with a series of lemmas. In what follows, an operator $S$ on $b\mathsf{X}$ is called globally stable when $S$ has a unique fixed point $\bar{v}$ in $b\mathsf{X}$ and $S^{n}v\to\bar{v}$ as $n\to\infty$ for all $v\in b\mathsf{X}$.
Lemma 5.6.
For each $\sigma\in\Sigma$, the operator $T_{\sigma}$ is globally stable on $b\mathsf{X}$.
Proof.
Fix $\sigma\in\Sigma$ and $v,w\in b\mathsf{X}$. Applying the triangle inequality, combined with the fact that $K$ is a positive linear operator, we have $|T_{\sigma}\,v-T_{\sigma}\,w|\leqslant|K(v-w)|\leqslant K|v-w|$. Since $\rho(K)<1$ and $b\mathsf{X}$ is a Banach lattice, by Proposition 4.1 and Theorem 2.1 in Stachurski and Zhang (2021), there exists some $N\in\mathbbm{N}$ such that $T_{\sigma}^{N}$ is a contraction. By Banach fixed point theorem, $T_{\sigma}^{N}$ is globally stable and it follows that $T_{\sigma}$ is globally stable on $b\mathsf{X}$. ∎
The next proof uses several concepts from Sargent and Stachurski (2025b). We refer to that paper for the definitions, which are omitted here for brevity.
Proof of Theorem 5.4.
To prove Theorem 5.4 we apply Theorem 5.5 of Sargent and Stachurski (2025b), which applies to the abstract dynamic program $(b\mathsf{X},\mathbbm{T})\coloneq(b\mathsf{X},\{T_{\sigma}:\sigma\in\Sigma\})$. First, in $b\mathsf{X}$ with its usual partial order, the statement $v_{n}\uparrow v$ is equivalent to $v_{n}(x)\uparrow v(x)$ in $\mathbbm{R}$ for each $x\in\mathsf{X}$. From this fact and the monotone convergence theorem, we obtain $Tv_{n}\uparrow Tv$ whenever $(v_{n})\subset b\mathsf{X}$ and $v_{n}\uparrow v\in b\mathsf{X}$. Thus, $(b\mathsf{X},\mathbbm{T})$ is order continuous. Also, $b\mathsf{X}$ is countably-Dedekind complete because the set of bounded Borel measurable functions is closed under pointwise convergence. In addition, each $T_{\sigma}$ is globally stable (Lemma 5.6) and hence order stable (see Sargent and Stachurski (2025b), Example 2.1). Finally, let $\hat{T}$ be the operator given by $\hat{T}v=|\pi|+Kv$. Since $\rho(K)<1$, there is a $u\in b\mathsf{X}_{+}$ with $\hat{T}u=u$. For this $u$ we have $T_{\sigma}\,u\leqslant\hat{T}u\leqslant u$, so $(b\mathsf{X},\mathbbm{T})$ is bounded above. It now follows from Theorem 5.5 of Sargent and Stachurski (2025b) that the value function is the unique solution to the Bellman equation and at least one optimal policy exists.
To prove the final claim we first note that $T$ is globally stable on $b\mathsf{X}$, since, given $v,w\in b\mathsf{X}$, we have $|Tv-Tw|\leqslant|K(v-w)|\leqslant K|v-w|$, so the argument in the proof of Lemma 5.6 applies. Now fix $v\in icb\mathsf{X}$. Since $K$ is invariant on $icb\mathsf{X}$, the element $-c+Kv$ is in $icb\mathsf{X}$. Moreover, $\pi\in icb\mathsf{X}$ and so $\max\{\pi,-c+Kv\}\in icb\mathsf{X}$. This shows that $T$ is invariant on $icb\mathsf{X}$. Since $icb\mathsf{X}$ is nonempty and closed under uniform limits, and $T$ is globally stable on $b\mathsf{X}$, we conclude that $v^{*}\in icb\mathsf{X}$. The proof of Theorem 5.4 is now done. ∎
Lemma 5.7.
If $v_{\sigma}(x)=v^{*}(x)$ for some $x\in\mathsf{X}$, then $v_{\sigma}=v^{*}$ almost everywhere.
Proof.
Since $v_{\sigma}\leqslant v^{*}$ and $T_{\sigma}\,v\leqslant Tv$ for all $v\in b\mathsf{X}$, we have $v_{\sigma}=T_{\sigma}\,v_{\sigma}\leqslant T_{\sigma}\,v^{*}\leqslant Tv^{*}=v^{*}$. As a result, $v_{\sigma}(x)=v^{*}(x)$ implies $(T_{\sigma}\,v_{\sigma})(x)=(T_{\sigma}\,v^{*})(x)$, which, using the definition of $T_{\sigma}$ in (16), yields
$\beta(x)\int v^{*}(x^{\prime})Q(x,\mathop{}\!\mathrm{d}x^{\prime})=\beta(x)\int v_{\sigma}(x^{\prime})Q(x,\mathop{}\!\mathrm{d}x^{\prime}).$
As $\beta(x)>0$, we obtain $\int(v^{*}(x^{\prime})-v_{\sigma}(x^{\prime}))Q(x,\mathop{}\!\mathrm{d}x^{\prime})=0$. If $v^{*}>v_{\sigma}$ on a set $E$ of positive Lebesgue measure, then, at the same time, we have
$\int(v^{*}(x^{\prime})-v_{\sigma}(x^{\prime}))Q(x,\mathop{}\!\mathrm{d}x^{\prime})\geqslant\int_{E}(v^{*}(x^{\prime})-v_{\sigma}(x^{\prime}))Q(x,\mathop{}\!\mathrm{d}x^{\prime})>0,$
where the last inequality is due to the assumption that $Q(x,\mathop{}\!\mathrm{d}x^{\prime})$ assigns positive measure to any such $E$. From this contradiction, we conclude that $v^{*}=v_{\sigma}$ almost everywhere. ∎
Lemma 5.8.
Let $f,g$ be two increasing functions in $b\mathsf{X}$ with $f\leqslant g$. If $g$ is continuous and $f=g$ almost everywhere, then $f=g$.
Proof.
If $f$ is continuous, the result follows from the fact that any two continuous measurable functions equal almost everywhere on an interval of $\mathbbm{R}$ are equal. Now suppose instead that $f$ is not continous at $x_{0}\in\mathsf{X}$. Let $a=\lim_{x\uparrow x_{0}}f(x)$ and $b=\lim_{x\downarrow x_{0}}f(x)$. Monotonicity and discontinuiy of $f$ implies $a<b$. Since $g$ is continuous, $g((a,b))^{-1}$ is open. We claim that $g((a,b))^{-1}$ is nonempty. Indeed, since $f=g$ almost everywhere, there are $y,y^{\prime}\in\mathsf{X}$ such that $g(y)\leqslant a$ and $g(y^{\prime})\geqslant b$. Since $g$ is continuous and increasing, there exist $y^{\prime\prime}\in\mathsf{X}$ such that $g(y^{\prime\prime})\in(a,b)$. Thus $g((a,b))^{-1}$ is a nonempty open set. This proves the existence of a set of nonzero Lebesgue measure on which $f<g$. Contradiction. ∎
Proof of Theorem 5.5.
Suppose there exists an $x\in\mathsf{X}$ and a policy $\sigma$ such that $v_{\sigma}(x)=v^{*}(x)$. Then, by Lemma 5.7, we have $v_{\sigma}=v^{*}$ almost everywhere. By Theorem 5.5, the function $v^{*}$ is increasing and continuous. By Lemma 5.6, the operator $T_{\sigma}$ is invariant on the closed set $ib\mathsf{X}$, so the fixed point is in $ib\mathsf{X}$; in particular, $v_{\sigma}$ is also increasing. By definition, $v_{\sigma}\leqslant v^{*}$. Hence, by Lemma 5.8, we have $v_{\sigma}=v^{*}$. ∎
5.4. A Two-State Example
In this section, we show that the irreducibility assumptions used in Theorem 3.3, Theorem 4.2, and Theorem 4.4 cannot be dropped: without irreducibility, (E1)(E3) are not generally equivalent. To make the example as simple as possible, we specialize to the case where $\mathsf{X}$ is finite and study the role of irreducibility in this setting. In applying these theorems, we take the metric on $\mathsf{X}$ to be the discrete metric, so that all subsets of $\mathsf{X}$ are open and $\mathscr{B}$ is all subsets of $\mathsf{X}$. We also assume that $\mathsf{A}$ is finite and impose the discrete metric on $\mathsf{A}$. As a result, every map and hence every policy from $\mathsf{X}$ to $\mathsf{A}$ is continuous.
Prior to stating our example, we note some elementary properties of irreducibility on discrete states. Given a transition kernel $Q$ on $\mathsf{X}$, we say that $y$ in $\mathsf{X}$ is $Q$-accessible from $x\in\mathsf{X}$ when there exists an $m\in\mathbbm{N}$ such that $Q^{m}(x,y)>0$. The usual definition of irreducibility of a transition kernel $Q$ on finite $\mathsf{X}$ is that, for every $x,y\in\mathsf{X}$, $x$ is $Q$-accessible from $y$ and $y$ is $Q$-accessible from $x$. To distinguish between different notions of irreducibility, we call this discrete irreducibility.
Lemma 5.9.
  1. (a)
    $P_{\sigma}$ is strongly irreducible.
  2. (b)
    $P_{\sigma}$ is discretely irreducible.
Proof.
((a) $\implies$ (b)) Fix $x,y\in\mathsf{X}$. Let $\mathbbm{1}_{x}(z)$ equal $1$ when $z=x$ and zero elsewhere. Let $\mathbbm{1}_{y}$ be defined analogously. Since $\mathbbm{1}_{x}\in b\mathsf{X}_{+}$ and $\mathbbm{1}_{y}\in(b\mathsf{X}_{+})^{\prime}$, there exists an $m\in\mathbbm{N}$ with $\left\langle{\mathbbm{1}_{x},P_{\sigma}^{m}\mathbbm{1}_{y}}\right\rangle>0$. But $\left\langle{\mathbbm{1}_{x},P_{\sigma}^{m}\mathbbm{1}_{y}}\right\rangle=P_{\sigma}^{m}(x,y)$, so $y$ is $P_{\sigma}$-accessible from $x$. This proves that $P_{\sigma}$ is discretely irreducible.
((b) $\implies$ (a)) Fix a nonzero $f\in b\mathsf{X}_{+}$ and nonzero $\mu\in(b\mathsf{X}_{+})^{\prime}$. Since $\mathsf{X}$ is finite, these element are just maps from $\mathsf{X}$ to $\mathbbm{R}_{+}$ and, as both are nonzero, we can find $\bar{x},\bar{y}\in\mathsf{X}$ such that $f(\bar{y})>0$ and $\mu(\bar{x})>0$. Moreover, since $P_{\sigma}$ is discretely irreducible, there is $m\in\mathbbm{N}$, such that $P_{\sigma}^{m}(\bar{x},\bar{y})>0$. As a result,
This proves that $P_{\sigma}$ is strongly irreducible. ∎
Now consider a two-state MDP with $\mathsf{X}=\{1,2\}$ and $\mathsf{A}=\{1,2\}$. The feasible coprrespondence is defined by $\Gamma(1)=\{1,2\}$ and $\Gamma(2)=\{2\}$. The reward function is defined by $r(i,j)=r_{ij}$ with
$P_{\sigma}=P_{\pi}=\begin{pmatrix}1&0\\ 0&1\end{pmatrix}.$
Observe that $P_{\sigma}=P_{\pi}$ fails to be discretely irreducible. By Lemma 4.3 and Lemma 5.9, this means that this transition kernel is neither strongly irreducible, nor weakly irreducible, nor open set irreducible.
Let us compute the lifetime value functions for the optimal $\sigma$ and $\pi$. For policy $\sigma$, we have $r_{\sigma}=(r_{12},r_{22})=(1,2)$ and $v_{\sigma}=(10,20)$. For $\pi$, we have $r_{\pi}=(r_{11},r_{22})=(0,2)$ and $v_{\pi}=(0,20)$. Since, $v^{*}(x)\coloneq\sup_{s\in\Sigma}v_{s}(x)=v_{\sigma}(x)$, we see that $\sigma$ is an optimal policy. On the other hand, $v_{\pi}(2)=v^{*}(2)$, but $v_{\pi}(1)<v^{*}(1)$, indicating that optimality in one state does not guarantee global optimality. This confirms that irreducibility cannot be dropped from the statements of Theorems 3.3, 4.2, and 4.4.
@@ 6. Extensions and Future Work
Using MDP optimality results from Bäuerle and Rieder (2011) or Bertsekas (2022), it should be possible to extend our results to the case of unbounded rewards by replacing the ordinary supremum norm on $b\mathsf{X}$ with a weighted supremum norm. Also, while our results have focused on standard MDPs with constant discount factors, one useful variation of this model is MDPs with state-dependent discount factors, so that $\beta$ becomes a map from $\mathsf{X}$ to $\mathbbm{R}_{+}$ (see, e.g., Stachurski and Zhang (2021)). We conjecture that the main results we obtained for MDPs will extend to generalized ADPs with state-dependent discounting under suitable stability and irreducibility assumptions. Finally, it seems likely that results similar to Theorem 3.3 will be valid for some continuous time MDPs, as well as at least some of the nonstandard discrete time dynamic programs discussed in Bertsekas (2022) and Sargent and Stachurski (2025a). These topics are also left for future work.